A government official is in charge of allocating social programs throughout the city of Vancouver. He will decide where these social outreach programs should be located based on the percentage of residents living below the poverty line in each region of the city. He takes a simple random sample of 129 people living in Gastown and finds that 25 have an annual income that is below the poverty line. The proportion of the 129 people who are living below the poverty line, 25/129
Use the sample data to compute a 95% confidence interval for the true proportion of Gastown residents living below the poverty line
_____,_____

Answer by Boreal(15235)

The 95% CI for a single proportion is mean +/- 1.96* sqrt{p*(1-p)/n}
sqrt term: p=25/129 or 0.194
sqrt 0.194*0.806/129=0.0348
the interval is 0.0348*1.96=0.068 in half width
(0.126, 0.262)